Circular No. 7714

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ed IN NGC 706 S. Benetti, G. Altavilla, E. Cappellaro, A. Pastorello, and M. Turatto,

Osservatorio Astronomico di Padova; and C. Contreras, J. Prieto, and A. Clocchiatti, Pontificia Universidad Catolica de Chile; and F. Patat, European Southern Observatory, report: “A noisy spectrum (range 380–900 nm; resolution 1 nm) of SN 2001ed (cf. IAUC 7703) was obtained on Sept. 9.257 UT using the Danish 1.54-m telescope (+ DFOSC) at La Silla. The spectrum shows the typical features of a type-Ia supernova having a phase of 22 ± 5 days. The expansion velocity deduced from the Si II 635.5-nm feature is ∼ 10 500 km/s, if we assume a recession velocity of 4980 km/s for the parent galaxy (NASA/IPAC Extragalactic Database).”

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001ed, obtained by P. Berlind on Sept. 11.42 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova several weeks after maximum light.

SUPERNOVAE 2001eb AND 2001ef Matheson et al. add that a spectrum of SN 2001eb (cf. IAUC 7702),

obtained by Berlind as above on Sept. 11.48 UT, shows it to be a type-Ia supernova several weeks after maximum light. A spectrum of SN 2001ef (cf. IAUC 7710), obtained by Berlind on Sept. 11.50, shows it to be a type-Ia supernova near maximum light. Adopting the NED recession velocity of 2476 km/s for the host galaxy, the supernova expansion velocity is ∼ 12 800 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days after maximum light.

SUPERNOVA 2001eh IN UGC 1162 M. Ganeshalingam and W. D. Li, University of California at Berkeley

(UCB), report the independent discovery of SN 2001eh (cf. IAUC 7712) by LOTOSS (cf. IAUC 7514) in unfiltered images taken with the Katzman Automatic Imaging Telescope. Available photometry (all unfiltered): Aug. 31.5 UT, [19.0; Sept. 4.5, 17.6; 5.5, 17.3; 7.5, 17.0; 9.5, 16.7; 10.5, 16.7. SN 2001eh is located 38′′.3 west and 4′′.7 north of the nucleus of UGC 1162. R. Chornock and A. V. Filippenko, UCB, report that a CCD spectrum (range 330–1000 nm) obtained on Sept. 11 with the 3-m Shane reflector at Lick Observatory reveals that SN 2001eh is a type-Ia supernova. The spectrum closely resembles that of SN 1990N, seven days before maximum brightness.

2001 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green